# Tilt your Head: Activating the Hidden Spatial-Invariance of Classifiers

### v1

@@ Abstract
Deep neural networks are applied in more and more areas of everyday life. However, they still lack essential abilities, such as robustly dealing with spatially transformed input signals. Approaches to mitigate this severe robustness issue are limited to two pathways: Either models are implicitly regularised by increased sample variability (data augmentation) or explicitly constrained by hard-coded inductive biases. The limiting factor of the former is the size of the data space, which renders sufficient sample coverage intractable. The latter is limited by the engineering effort required to develop such inductive biases for every possible scenario. Instead, we take inspiration from human behaviour, where percepts are modified by mental or physical actions during inference. We propose a novel technique to emulate such an inference process for neural nets. This is achieved by traversing a sparsified inverse transformation tree during inference using parallel energy-based evaluations. Our proposed inference algorithm, called Inverse Transformation Search (ITS), is model-agnostic and equips the model with zero-shot pseudo-invariance to spatially transformed inputs. We evaluated our method on several benchmark datasets, including a synthesised ImageNet test set. ITS outperforms the utilised baselines on all zero-shot test scenarios.
Canonicalisation, Geometric Deep Learning
@@ 1 Introduction
A primary objective in Deep learning is to develop models that exhibit robust generalisation during inference (commonly evaluated using test sets). To improve generalisation, models are trained with data sets of increasing variability and size. However, this often requires using larger models (Chefer et al., 2022). As novel datasets more closely mirror real-world scenarios, models undergo rigorous assessments in diverse, dynamic environments. This requires the models to learn intricate functions that generalise to all the variations presented in the data. Each new variation (dimension) scales the data space exponentially, known as the combinatorial explosion. Sampling in such spaces gets progressively harder (curse of dimensionality). For instance, consider an object from infinitely many perspectives. In a standard image classification task, all these variations have the same label assigned. Using 2D spatial transformations, these variations bring only redundant information. Hence, considering every conceivable variant is impractical, computationally infeasible, and even useless in terms of information gain. If a model can correctly predict an object from all perspectives, it is (pseudo-)invariant to these perspective changes.
@@
Inductive biases for spatial symmetries are commonly developed for CNNs by exploiting their weight-shared filters (Schmidt & Stober, 2022). Instead of regularising the filter space using data, it can be constrained explicitly to symmetric weight matrices (so-called isotropic filters) (Hu et al., 2019). Invariance can also be achieved without such constraints by considering all transformed filter variations followed by a max-out (Gens & Domingos, 2014; Cohen & Welling, 2016; Bao & Song, 2020). This can be formulated more flexibly by a steerable basis using Bessel functions (Qiu et al., 2018; Delchevalerie et al., 2021) or 2D Hermite polynomials to allow for weight sharing across kernel orientations (Ecker et al., 2019; Ustyuzhaninov et al., 2020). However, these methods are limited to specific transformations, like centre rotations. Zhou et al. (2020) proposed to meta-learn the filter structure (binary mask) based on the equivariance required to solve a task.
Unfortunately, these inductive biases come with a significant engineering effort. The same applies to methods, which learn the geometry of the latent space (Arvanitidis et al., 2017) to find equivariance paths. With the introduction of Spatial Transformer Network (STN) (Jaderberg et al., 2015), another promising avenue began. These networks remove the spatial perturbations, i.e., canonicalise the transformed input signals by learning a highly non-linear mapping from pixel to transformation space. A more effective alternative is the mapping to canonical coordinate spaces (Esteves et al., 2018; Tai et al., 2019). To learn transformation compositions, STNs can be stacked (Lin & Lucey, 2017; Detlefsen et al., 2018). These transformations can also act on each pixel individually using pixel-wise motion fields (Shu et al., 2018). However, all of these variants require augmented training data. We propose a novel inference strategy that canonicalises inputs without augmented training.
@@ 2 Preliminaries
Understanding how to undo transformations requires a basic comprehension of symmetries and their mathematical formulations as groups. A symmetry of an object is a transformation that preserves the object’s structure. For instance, rotating an equilateral triangle around its centre by $0^{\circ}$, $120^{\circ}$ or $240^{\circ}$ leaves the triangle unaffected, as illustrated in Figure 2. These rotations form a group of finite size, i.e., a finite group.
@@
Let us extend the picture from discrete to continuous structures. For a circle, centre rotations of any angle send the object back to itself. Formally, these groups are called Matrix Lie groups (Hall, 2015), which form a smooth differentiable manifold. For the basic comprehension of this work, it is only important to understand that group elements can be represented by parameterised matrices. In this work, we focus on centre rotation, isotropic scaling, and shearing transformations (and their compositions). All of them are represented by $2\times 2$ matrices $\mathbf{T}$. This particular group is called the Special Linear Group $\text{SL}_{2}(\mathbb{R})$ (Hall, 2015).
@@
Given $g\in G$, $g$ can be used as a generator to form a finite (cyclic) subgroup $\{\langle g\rangle^{k}\mid k\in\mathbb{Z}\}$. $\langle g\rangle^{k}$ denotes $g$ to be a generator acting $k$-times.
@@ 3 Localizing Canonical Forms
We derive an energy-based evaluation from the concept of invariances. We assume that the function $f_{\mathbf{\theta}}$ is sensitive to $G$, which is true for all neural networks when $G$ is getting more complex. That is, for any non-trivial element $g$ in $G$ we have $f_{\mathbf{\theta}}(\mathbf{x})\neq f_{\mathbf{\theta}}(g\mathbf{x})$. However, based on Def. 2.1, there exists a $g^{-1}\in G$ such that $f_{\mathbf{\theta}}(\mathbf{x})=f_{\mathbf{\theta}}(g^{-1}g\mathbf{x})$. Using a confidence measure $\rho:f_{\mathbf{\theta}}(\cdot)\to\mathbb{R}$ (Kaba et al., 2023), we can determine $g^{-1}$ by
@@
The confidence measure should concentrate all its mass around the canonical form. However, we found that both baselines tend to increase their mass towards the domain borders. This is best observed in the rotation experiment in Figure 3 (last row). Here, the spatial alignment bias of FashionMNIST (Xiao et al., 2017) leads to increased confidence for $90^{\circ}$ rotated images (high activation of horizontal edge detectors). Similar confidence bumps appear when scaling up or down a sample until pixelation or disappearance. We propose a more stable measure, which does not suffer from these confidence anomalies.
@@
Without explicit training or inductive biases, neural nets corrupt the geometry of group orbits during encoding (Arvanitidis et al., 2017). When traversing the orbit’s trajectory, the energy will fluctuate (Li et al., 2018). As both confidence scores utilise the energy, the confidence will fluctuate likewise. We remove noise from the energy by estimating its expected value $\bar{E}_{\mathbf{\theta}}(\cdot)$ using Monte-Carlo Dropout. This renders the measure stochastic. We use a generator $\hat{g}$ to generate the finite group $G$. We can evolve $\mathbf{x}$ along $\text{orb}_{G}(\mathbf{x})$ by incrementing or decrementing $n$ in $\hat{g}^{n}\mathbf{x}$.
Instead of evaluating each point independently, we consider the behaviour in the local neighbourhood as well. We use the curvature of the energy along the orbit. Local maxima have high negative curvature, which we approximate by the second derivative. This leads to the confidence measure,
@@
where $K$ is a Gaussian smoothing kernel to reduce local over-confidence. To ensure that the confidence decreases at domain borders, we use “nearest” padding. In practice, we can approximate Equation 7 as follows.
@@
Find the proofs in Section D.1 and Section D.2.
@@ 4 Inverse Transformation Search
Figure 4: Our proposed Inverse Transformation Search (ITS) on a perturbed image $\hat{\mathbf{x}}$. In each level of the search tree, the orbit of one $G$ is evaluated. At each level all samples are evaluated using Equation 7 and a pre-trained model $f_{\theta}$.
Figure 5: An example iteration of our algorithm with a generated sub-group of size $9$ and $2$ hypotheses. The first layer rotates the input, the second layer scales it, and the last layer shears it.
An overview and an example of our proposed search procedure are illustrated in Figure 4. In the following, we spell out the most important aspects of it with more details in Appendix C.
@@
We assume that a vanilla data point $\mathbf{x}$ is transformed by some groups acting on it. Let $\hat{\mathbf{x}}:=(g_{0}\circ g_{1}\circ\ldots\circ g_{m})\mathbf{x}$ where $g_{0},g_{1},\ldots,g_{m}$ might be elements of different groups. We follow a heuristic (greedy) approach by decomposing the transformation into different groups acting sequentially on the input. This reduces the complexity of the problem significantly. We propose an iterative search procedure, which aims to eliminate one $g_{k}$ at every step using its inverse $g_{k}^{-1}$. This assumes either that the order of transformation is known, which is unlikely in real-world settings or that we deal with Abelian groups. We show that the restrictions imposed by the latter can be attenuated, as only a sub-sequence needs to be Abelian.
@@
The proof can be found in Section D.3. To cope with Lie groups, we generate sub-groups of size $|G|$ using generators. In that light, each layer comprises $|G|$ elements (as a finite set of equidistant samples along the orbit). This allows us to express transformations by an exponential $\hat{g}^{n}$ with $n\in\mathbb{N}$. This renders the evaluation in each layer tractable, as only a finite set of points needs to be evaluated. [^1] All non-cyclic groups, like translations, would at some point shift the image signal outside the visible pixel grid. In practice, we consider only a reasonably bounded set of transformations.
@@
In theory, such a dynamic search process over multiple hypotheses would allow the model to ”change its mind”. That is, rejecting $H_{0}$ (the initially dominating hypothesis) and instead accepting some other $H$ for the final incumbent (output of the inference process). Figure 5 already provides an example for such phenomena, which we investigate more thoroughly using Figure 6. We constructed a regular grid of rotated and sheared Fashion-MNIST test sets. Each point in Figure 6 is the average performance of a CNN with ITS (see Section B.2 for more details). Colour changes from layers 1 to 2 indicate that the initial hypothesis is on average superseded by the other. To maximise the performance outcome (we focus on classification accuracy), we need to substitute the leading hypothesis $H_{0}$ with its alternative $H_{1}$ in some regions. Without a ground-truth signal (during test time), we need a measure to determine when to swap hypotheses. In our experiments, we used class support, which is the number of occurrences of the predicted class over the entire tree. However, this indicator has only minor advantages over the confidence score (see Figure 6 e,f). We leave this promising avenue for future research.
@@
We have a fixed-width $k$-ary tree, where the number of elements in the first level is $n:=|G|$ and in all consecutive levels $kn$. As all evaluations are performed in parallel, $n$ does not affect the runtime. For our implementation, the Monte Carlo estimates are performed sequentially scaling the complexity by $M$. This gives $\mathcal{O}(Mk^{d})$, where $d$ denotes the depth of the tree. We quantified the runtime cost in Section B.1. Our inference strategy is particularly interesting for applications where it is worth trading runtime against robustness (Ghamizi et al., 2022).
@@ 5 Experiments
If not further specified, we used zero-padding to define areas outside the pixel space $\Omega$, bilinear interpolation, and a group cardinality of $n=17$. We used two CNN-backbones simple as the LeNet architecture (Lecun et al., 1998) (see Section A.1). All models are trained with the AdamW optimizer (Loshchilov & Hutter, 2017) using default parameters. We minimised the negative log-likelihood using ground-truth image labels. We used a learning rate of $35e-$ 3, $3$ epochs for MNIST, $5$ epochs for Fashion-MNIST and mini-batches of size $128$. In all MNIST experiments, we removed the class $9$ samples, as this introduces a class-superposition for $180^{\circ}$ rotations. You can find more details in our publicly available source code. [^2] Added after blind reviews.

### v2

@@ Abstract
Deep neural networks are applied in more and more areas of everyday life. However, they still lack essential abilities, such as robustly dealing with spatially transformed input signals. Approaches to mitigate this severe robustness issue are limited to two pathways: Either models are implicitly regularised by increased sample variability (data augmentation) or explicitly constrained by hard-coded inductive biases. The limiting factor of the former is the size of the data space, which renders sufficient sample coverage intractable. The latter is limited by the engineering effort required to develop such inductive biases for every possible scenario. Instead, we take inspiration from human behaviour, where percepts are modified by mental or physical actions during inference. We propose a novel technique to emulate such an inference process. This is achieved by traversing a sparsified inverse transformation tree during inference using parallel energy-based evaluations. Our proposed inference algorithm, called Inverse Transformation Search (ITS), is model-agnostic and equips the model with zero-shot pseudo-invariance to spatially transformed inputs. We evaluated our method on several benchmark datasets, including a synthesised ImageNet test set. ITS outperforms the utilised baselines on all zero-shot test scenarios.
Canonicalisation, Geometric Deep Learning, Computer Vision
@@ 1 Introduction
A primary objective in Deep learning is to develop models that exhibit robust generalisation during inference (commonly evaluated using test sets). To improve generalisation, models are trained with data sets of increasing variability and size. However, this often requires using larger models (Chefer et al., 2022). As novel datasets more closely mirror real-world scenarios, models undergo rigorous assessments in diverse, dynamic environments. This requires the models to learn intricate functions that generalise to all the variations presented in the data. It was shown that modern vision models break even under minor input perturbations (Engstrom et al., 2019), posing severe risks for real-world applications. Each new variation (dimension) scales the data space exponentially, known as the combinatorial explosion. Sampling in such spaces gets progressively harder (curse of dimensionality). For instance, consider an object from infinitely many perspectives. In a standard image classification task, all these variations have the same label assigned. Using 2D spatial transformations, these variations bring only redundant information. Hence, considering every conceivable variant is impractical, computationally infeasible, and even useless in terms of information gain. If a model can correctly predict an object from all perspectives, it is (pseudo-)invariant to these perspective changes.
@@
Inductive biases for spatial symmetries are commonly developed for CNNs by exploiting their weight-shared filters (Schmidt & Stober, 2022). Instead of regularising the filter space using data, it can be constrained explicitly to symmetric weight matrices (so-called isotropic filters) (Hu et al., 2019). Invariance can also be achieved without such constraints by considering all transformed filter variations followed by a max-out (Gens & Domingos, 2014; Cohen & Welling, 2016; Bao & Song, 2020). This can be formulated more flexibly by a steerable basis for the space of convolutional kernels (Weiler et al., 2017). This allows for weight sharing across filter orientation without aliasing artifacts. Recently, various options for such basis functions were proposed, most of which are derived from circular symmetries. This includes Bessel functions (Qiu et al., 2018; Delchevalerie et al., 2021), Hermite polynomials (Ecker et al., 2019; Ustyuzhaninov et al., 2020) and irreducible representations to steer the harmonics (Weiler & Cesa, 2019). Zhou et al. (2020) proposed to meta-learn the filter structure (binary mask) based on the equivariance required to solve a task. Recently, it was shown that equivariant models can be canonicalised by fitting a prior over the group during training (Kaba et al., 2023; Mondal et al., 2023).
Unfortunately, these inductive biases come with a significant engineering effort, which renders a generalisation to other symmetries cumbersome. The same applies to methods, which learn the geometry of the latent space (Arvanitidis et al., 2017) to find equivariance paths. With the introduction of Spatial Transformer Network (STN) (Jaderberg et al., 2015), another promising avenue began. These networks remove the spatial perturbations, i.e., canonicalise the transformed input signals by learning a highly non-linear mapping from pixel to transformation space. A more effective alternative is the mapping to canonical coordinate spaces (Esteves et al., 2018; Tai et al., 2019). To learn transformation compositions, STNs can be stacked (Lin & Lucey, 2017; Detlefsen et al., 2018). These transformations can also act on each pixel individually using pixel-wise motion fields (Shu et al., 2018). However, all of these variants require augmented training data. We propose a novel inference strategy that canonicalises inputs without augmented training.
@@ 2 Preliminaries
Understanding how to undo transformations requires a basic comprehension of symmetries and their mathematical formulations as groups. A symmetry of an object is a transformation that preserves the object’s structure. For instance, rotating an equilateral triangle around its centre by $0^{\circ}$, $120^{\circ}$ or $240^{\circ}$ leaves the triangle unaffected, as illustrated in Figure 2 (left). These rotations form a group of finite size, i.e., a finite group.
@@
Let us extend the picture from discrete to continuous structures. For a circle, centre rotations of any angle send the object back to itself. Formally, these groups are called Matrix Lie groups (Hall, 2015), which form a smooth differentiable manifold. For the basic comprehension of this work, it is only important to understand that group elements can be represented by parameterised matrices. In this work, we focus on centre-rotation, isotropic-scaling, and shearing transformations (and their compositions). All of them are represented by $2\times 2$ matrices $\mathbf{T}$. This particular group is called the Special Linear Group $\text{SL}_{2}(\mathbb{R})$ (Hall, 2015).
@@
Given $g\in G$, $g$ can be used as a generator to form a finite (cyclic) subgroup $\{\langle g\rangle^{k}\mid k\in\mathbb{Z}\}$. $\langle g\rangle^{k}$ denotes $g$ to be a generator acting $k$-times. To remove visual clutter, we use $\hat{g}$ to denote the generating group element.
@@ 3 Localizing Canonical Forms
We derive an energy-based evaluation from the concept of invariances. We assume that the function $f_{\mathbf{\theta}}$ is sensitive to $G$, which is true for all neural networks when $G$ is getting more complex. That is, for any non-trivial element $g$ in $G$ we have $f_{\mathbf{\theta}}(\mathbf{x})\neq f_{\mathbf{\theta}}(g\mathbf{x})$. However, based on Def. 2.1, there exists a $g^{-1}\in G$ such that $f_{\mathbf{\theta}}(\mathbf{x})=f_{\mathbf{\theta}}(g^{-1}g\mathbf{x})$. Inspired by Kaba et al. (2023) we can determine $g^{-1}$ using a confidence measure $\rho:f_{\mathbf{\theta}}(\cdot)\to\mathbb{R}$, such that
@@
The confidence measure should concentrate all its mass around the canonical form. However, we found that both baselines tend to increase their mass towards the domain borders. This is best observed in the rotation experiment in Figure 3 (last row). Here, the spatial alignment bias of FashionMNIST (Xiao et al., 2017) leads to increased confidence for $90^{\circ}$ rotated images (high activation of horizontal edge detectors). Similar confidence bumps appear when scaling up or down a sample until pixelation or disappearance. Without explicit training or inductive biases, neural networks corrupt the geometry of group orbits during encoding (Arvanitidis et al., 2017). When traversing the orbit’s trajectory, the energy will fluctuate (Li et al., 2018). As both confidence scores utilise the energy, the confidence will fluctuate likewise.
@@
We propose a more stable measure, which does not suffer from these confidence anomalies. First, we remove noise from the energy by estimating its expected value $\bar{E}_{\mathbf{\theta}}(\cdot)$ using Monte-Carlo Dropout. This renders the measure stochastic (as the Bayesian-induced confidence), reducing the impact of outliers. To further mitigate instabilities, we evaluate local neighbourhoods instead of single point estimates. This can be achieved by convolving the energy estimate over the orbit with a kernel. Lastly, we compute the negative curvature of the resulting energy surface instead of using the energy estimate directly. This allows us to reduce the confidence mass at the domain borders by using “nearest” padding, which lowers the curvature at these points. We approximate the curvature by the second derivative, such that
@@
where $K$ is a Gaussian smoothing kernel to reduce local over-confidence. We use a generator $\hat{g}$ to generate the finite group $G$. Therefore, we can evolve $\mathbf{x}$ along $\text{orb}_{G}(\mathbf{x})$ by incrementing or decrementing $n$ in $\hat{g}^{n}\mathbf{x}$. In practice, we can approximate Equation 7 as follows.
@@
Find the proofs in Section E.1 and Section E.2.
A comparison of all three measures is depicted in Figure 3. Our group-induced confidence either surpasses both baselines or achieves on par results.
@@ 4 Inverse Transformation Search
Figure 4: Our proposed Inverse Transformation Search (ITS) on a perturbed image. In each level of the search tree, the orbit of one subgroup is evaluated. All samples are evaluated using Equation 7 and a pre-trained classifier.
Figure 5: An example iteration of our algorithm with two hypotheses $H_{0}$ and $H_{1}$, and three generated subgroups, each with a cardinality of $|G|=9$. The first layer rotates the input, the second layer scales it, and the last layer shears it.
An overview and an example of our proposed search procedure are illustrated in Figure 4. In the following, we spell out the most important aspects of it with more details in Appendix A and Appendix D (including pseudocode).
@@
We deal with problem scenarios, where samples are perturbed by a some unknown affine transformation $g$ during test-time. We assume that $g$ decomposes into various subgroup transformations, $g=g_{0}\circ g_{1}\circ\ldots\circ g_{m}$. We propose an iterative search procedure, which aims to eliminate one $g_{k}$ at every step using its inverse $g_{k}^{-1}$. This heuristic (greedy) approach reduces the complexity of the problem significantly. However, this comes with the assumption that either the order of transformation is known, which is unlikely in real-world settings, or that we deal with Abelian groups. We show that the restrictions imposed by the latter can be attenuated, as only a sub-sequence needs to be Abelian.
@@
The proof can be found in Section E.3.
To cope with Lie groups, we generate subgroups of size $|G|$ using generators as already used in our group-induced confidence measure (see Theorem 3.1). In that light, each layer comprises $|G|$ elements (as a finite set of equidistant samples along the orbit). This allows us to express transformations by an exponential $\hat{g}^{n}$ with $n\in\mathbb{N}$. This renders the evaluation in each layer tractable, as only a finite set of points needs to be evaluated. [^1] All non-cyclic groups, like translations, would at some point shift the image signal outside the visible pixel grid. In practice, we consider only a reasonably bounded set of transformations.
@@
In theory, such a dynamic search process over multiple hypotheses would allow the model to ”change its mind”. That is, rejecting $H_{0}$ (the initially dominating hypothesis) and instead accepting some other $H$ for the final incumbent (output of the inference process). Figure 5 already provides an example for such phenomena, which we investigate more thoroughly in Figure 6. We constructed a regular grid of rotated and sheared Fashion-MNIST test sets. Each point in Figure 6 is the average performance of a CNN with ITS (see Section C.3 for more details). Colour changes from layers 1 to 2 indicate that the initial hypothesis is on average superseded by the other. To maximise the performance outcome (we focus on classification accuracy), we need to substitute the leading hypothesis $H_{0}$ with its alternative $H_{1}$ in some regions. Without a ground-truth signal (during test time), we need a measure to determine when to swap hypotheses. In our experiments, we used class support, which is the number of occurrences of the predicted class over the entire tree. However, this indicator has only minor advantages over the confidence score (see Figure 6 e,f). We leave this promising avenue for future research.
@@
We have a fixed-width $k$-ary tree, where the number of elements in the first level is $n:=|G|$ and in all consecutive levels $kn$. As all evaluations are performed in parallel, $n$ does not affect the runtime. For our implementation, the Monte Carlo estimates are performed sequentially, scaling the complexity by $M$. This gives $\mathcal{O}(Mk^{d})$, where $d$ denotes the depth of the tree. We quantified the runtime cost in Section C.1.
Application Domains
Given the runtime costs, we argue that our inference strategy is particularly interesting for applications where it is worth trading runtime against robustness. This includes medical applications, like X-ray classification (Ghamizi et al., 2022); analysis of satellite or drone imagery, like land suitability prediction for agriculture (Taghizadeh-Mehrjardi et al., 2020), monitoring crop health (Inoue, 2020), or identifying deforestation (Andrade et al., 2020); classifying celestial objects in telescopic images (Gómez et al., 2020) or analysing protein structures in biology from microscopic images (Moen et al., 2019).
@@ 5 Experiments
If not further specified, we used zero-padding to define areas outside the pixel space $\Omega$, bilinear interpolation, and a group cardinality of $n=17$. For our experiments on (Fahsion)-MNIST and GTSRB we used two CNN-backbones simple as the LeNet architecture (Lecun et al., 1998) (see Section A.2). These models are trained with the AdamW optimizer (Loshchilov & Hutter, 2017) using default parameters. We minimised the negative log-likelihood using ground-truth image labels. We used a learning rate of $35e-$ 3, $3$ epochs for MNIST, $5$ epochs for Fashion-MNIST, $10$ epochs for GTSRB and mini-batches of size $128$. In all MNIST experiments, we removed the class $9$ samples, as this introduces a class-superposition for $180^{\circ}$ rotations. More details can be found in our publicly available source code.222www.github.com/johSchm/ITS
@@ 5.2 Zero-Shot Affine Canonicalisation
Similar to STNs, we can undo affine transformations (i.e., canonicalize inputs). We compare ITS to a STN (Jaderberg et al., 2015) with one and three transformation layers using inverse compositional layers (Lin & Lucey, 2017), a Diffeomorphic Transformer Network (DTN) (Detlefsen et al., 2018), and a Equivariant Transformer Network (ETN) (Tai et al., 2019). To study the impact of depth, we contrasted a $3$- and a $5$-layer ITS. We used a layer stack of rotation + scaling + shearing + (rotation + scaling). We refer to these as ITS 3 and ITS 5, respectively. Table 1 shows the quantitative results of our experiments. The associated qualitative results as reconstructions (after applying the predicted inverse transformation) can be found in Figure 8.
We found all baselines incapable of performing affine canonicalisations outside the training data distribution. This makes sense, as their localisation networks (regressing the transformation parameters) learn pixel-to-transformation mappings from training data. They are unable to generalise - particularly to rotations, which are mostly present during testing. The STN and the DTN estimate the composed transformation matrix in one forward pass. They need to extrapolate from nearly no rotations during training to $360^{\circ}$ rotations during testing. As the rotation manifold is isomorphic to the unit circle ($\cos(\theta),\sin(\theta)$), this non-linearity needs to be learned implicitly. The ETN on the other hand uses canonical coordinates, which linearise the relationship between rotations.
5.3 Scaling up to ImageNet